find the extreme values of the function f(x,y,z)=x^(2)+2y^(2 )subject to the constraint x^(2)+y^(2)-z^(2)=1
Answer #1

Put x=1=, y=0 (in equation ) Solution = (d=1 So that value of 2-0 always in +(59,2) ton maxinum valm of & (892) x² + y = 1 +Point (0,0,0), (1,2,0), (1,1,0) +0,0,0) = 0 3 + (1,2,0)= + 2x2² = 9 +(1,1,0) = 12 + 2x2 = 3 So that the maximum valu of t(X,9